OpenStudy (anonymous):

If 3+3i and z2=7(cos(5pi/9)+isin(5pi/9)), then z1z2=____? A. 21sqrt2(cos(5pi/36)+isin(5pi/36)) B. (3sqrt2/7)(cos(-11pi/36)+isin(-11pi/36)) C. 21(cos(5pi/36)+isin(5pi/36)) D. 21sqrt2(cos(29pi/36)+isin(29pi/36)) I think it's supposed to say if Z1=3+3i, but I'm not entirely sure. How do?

jimthompson5910 (jim_thompson5910):

`I think it's supposed to say if Z1=3+3i` sounds like it

jimthompson5910 (jim_thompson5910):

z1 = 3+3i is in the form a+bi a = 3 b = 3 r = sqrt(a^2 + b^2) = ??? theta = arctan(b/a) = ???

OpenStudy (anonymous):

Ran out of time, the answer was D.